# Prospects for Cosmological Research with the FAST Array: 21-cm Intensity Mapping Survey Observation Strategies

### v1

@@ II.3 Foreground
II.4 Instrument parameters
In this paper, we consider several different configurations of FAST and Square Kilometre Array Phase 1 - Mid frequency (SKA1-Mid) for simulating the 21 cm IM. The specific instrument parameters are listed in Table 2.
The current capability of FAST to detect 21 cm signals covers the redshift range $0.350<z<$ 0.35. A new wide-band receiver under test extends this range to $1.840<z<$ 1.84 Zhang et al. (2023b). Previous galaxy surveys, however, have performed precise BAO measurements within $z<2.5$, achieving 1%-2% accuracy Alam et al. (2021); Aubourg et al. (2015). This paper assumes the upgraded new wide-band receiver on FAST can detect 21 cm signals up to $2.50<z<$ 2.5. Additionally, constructing 6$\times$40 m dishes or 20$\times$40 m dishes around FAST is considered. These dishes will use the same receiver and feed system as FAST, so their frequency bands and $N_{\mathrm{b}}$ will be identical. Given the outstanding performance of FAST, we also consider constructing five additional identical FAST telescopes to form FASTA, significantly improving survey efficiency. Furthermore, we consider equipping FAST with a PAF to increase the value of $N_{\mathrm{b}}$ to 100.
Regarding instrument temperature, in Ref. Jiang et al. (2020) the author measured the temperature of FAST near boresight to be 20 K in the L-band. We assume that the frequency band can be broadened while maintaining the instrument temperature at 20 K. For the PAF, we assume the temperature to be approximately 30 K due to the increase in the number of beams. The total observation time is assumed to be 10000 hours. The calculation for the survey area is as follows
$S_{\text{area}}=\eta\int_{\phi}\int_{\theta}\sin\theta\mathrm{d}\theta\mathrm{d}\phi\ ,$ (7)
where $\eta$ represents the fraction of the initial survey area that remains usable for scientific observation after excluding the Milky Way. For FAST/FASTA, we considered $\eta=0.7$ and the range of declination from $-14^{\circ}$ to $+66^{\circ}$, yielding an area of approximately 8100 $\text{deg}^{2}$.
For SKA1-Mid, we considered band I ($30.35<z<$ 3) planned for 21 cm IM, with the survey area of 20000 deg2 and the total observation time of 10000 hours. The instrument temperature is as follows Bacon et al. (2020)
$T_{\mathrm{inst}}=15\mathrm{~{}K}+30\mathrm{~{}K}\left(\frac{f}{\mathrm{GHz}}-0.75\right)^{2}.$ (8)
II.5 The Fisher matrix
@@
$C^{\mathrm{tot}}=C^{\mathrm{S}}+C^{\mathrm{N}}+C^{\mathrm{F}}.$ (9)
For the combination of different telescopes, we consider two scenarios. The first scenario involves two telescopes in close geographic proximity observing the same area of the sky, such as combining 40 m dishes and FAST. This approach maximizes the reduction of thermal noise impact on observations. Since $C^{\mathrm{S}}$ and $C^{\mathrm{F}}$ are only related to the observed sky area, they remain fixed. Different telescopes have varying $C^{\mathrm{N}}$ due to their differing thermal noise integration effects. Therefore, the total covariance matrix can be expressed as
$C^{\mathrm{tot}}=C^{\mathrm{S}}+\left((C_{1}^{\mathrm{N}})^{-1}+(C_{2}^{\mathrm{N}})^{-1}\right)^{-1}+C^{\mathrm{F}},$ (10)
@@
$C^{\mathrm{tot}}=\left((C_{1}^{\mathrm{tot}})^{-1}+(C_{2}^{\mathrm{tot}})^{-1}\right)^{-1},$ (11)
where $C_{1}^{\mathrm{tot}}$ and $C_{2}^{\mathrm{tot}}$ are the total covariance matrices of the different dishes.
After obtaining the total Fisher matrix $C^{\mathrm{tot}}$, we can transform it into the Fisher matrix of observations $\{p_{i}\}$ Bull et al. (2015)
$F_{ij}=\frac{1}{2}U_{\text{bin}}\int\frac{\mathrm{d}^{2}q\,\mathrm{d}y}{(2\pi)^{3}}\left[\frac{\partial\ln C^{\mathrm{tot}}(\bm{q},y)}{\partial p_{i}}\frac{\partial\ln C^{\mathrm{tot}}(\bm{q},y)}{\partial p_{j}}\right].$ (12)
In this paper, we mainly calculate the Fisher matrix for several key observables to constain the evolution of LSS, as well as the growth and geometry of the universe. These include the transverse distance indicator $D_{\mathrm{A}}(z)$, radial distance indicator $H(z)$ and the parameter combination $f\sigma_{8}(z)$ from redshift space distortions (RSDs), where $f(z)$ is growth rate and $\sigma_{8}$ is the amplitude of matter density fluctuations on a scale of $8h^{-1}$ Mpc (with $h$ the dimensionless Hubble constant).
@@
Figure 2: Relative errors of $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ for 21 cm IM with FAST, 40 m dishes, FAST+$6\times 40$ m dishes, and FAST+$20\times 40$ m dishes.
II.6 Cosmological parameter constraints
We use the Fisher matrix to propagate errors from observational errors to the BAO parameters $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$. Then, we use the errors of BAO parameters to form the likelihood function for cosmological models. Finally, we employ Markov Chain Monte Carlo (MCMC) analysis to constrain the cosmological parameters Witzemann et al. (2018). To compare with the mainstream observations, we calculated the constraints from the latest observations. The data used include: CMB – Planck 2018 $-\ell$ TT, TE, EE, and $-\ell$ TT, EE, and PR4 lensing data Aghanim et al. (2020); Efstathiou and Gratton (2021); Carron et al. (2022a, b); Rosenberg et al. (2022); BAO – DESI data Adame et al. (2024a); and supernova (SN) – Pantheon+ data Scolnic et al. (2022). The data combination of CMB, BAO, and SN in this paper is also abbreviated as CBS.
The equation of state (EoS) of dark energy is given by $w=p/\rho$, where $p$ is the pressure and $\rho$ is the energy density. We employ different models to describe its dynamics. (i) $\Lambda$CDM: $w=-1$. (ii) $w$CDM: $w(z)=w$, where $w$ is a constant EoS for dark energy. (iii) $w_{0}w_{a}$: $w$ varies with redshift according to the parameterization $w(z)=w_{0}+w_{a}z/(1+z)$, where $w_{0}$ is the present-day value of the equation of state parameter and $w_{a}$ describes its evolutionary behavior.
III Results and Discussion
@@
Figure 4: Constraints ($1\sigma$ and $2\sigma$ confidence level) on the dark energy models using simulated data of 21 cm IM with SKA1-Mid, FAST, FASTA, and observational data of CBS. Left, middle, and right panels show the constraints on the $\Lambda$CDM, $w$CDM, and $w_{0}w_{a}$ models, respectively.
@@
III.1 Combination of FAST and SKA-Mid
From Figure 1, it is evident that FAST and SKA1-Mid have distinct advantages across different redshift ranges. SKA1-Mid achieves tight constraints within the range of $0.80.35<z<$ 0.8 due to its high survey efficiency. In this range, the relative errors in BAO parameters are typically below 3%. On the other hand, FAST attains robust constraints within the range of $2.50.8<z<$ 2.5 because of its high transverse resolution and sensitivity. In this range, the relative errors in BAO parameters are typically around or below 4%. Their combined observations enable precise measurement of BAO parameters across the redshift range of $2.50<z<$ 2.5, achieving constraints on $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ to below 2% for most of this range, and to as low as 1% at $z\approx 1$.
III.2 Combination of FAST and 40 m dishes
In this subsection, we discuss the extent to which incorporating 40 m dishes with FAST enhances the constraints on BAO parameters in 21 cm IM. An array of 40 m dishes, located in close geographical proximity to FAST, enhances the survey efficiency by enabling observations of the same sky regions.
Figure 2 illustrates the fractional constraints achievable on $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ using 21 cm IM with an array of 40 m dishes, FAST, and their combined configurations. The 40 m dishes perform better at low redshifts because they have higher survey efficiency. FAST is more effective at higher redshifts due to its greater sensitivity and resolution. The combination of FAST and 40 m dishes significantly reduces observational errors, particularly at high redshifts. The relative errors in BAO parameters are typically below 4%, with errors reaching as low as 2% at $z\approx 1$.
For the configurations of $6\times 40$ m dishes and $20\times 40$ m dishes combined with FAST, Figure 2 shows that the larger array provides only marginal improvements. This result suggests that the combination of FAST with $6\times 40$ m dishes offers a more optimal configuration for the 21 cm IM survey.
@@ III.2 Combination of FAST and 40 m dishes
III.3 FASTA
[…]
Compared to the results in Ref. Wu and Zhang (2022) for FAST ($0.350<z<$ 0.35), a wide-band receiver proposed in this paper utilizes the resolution advantage of the FAST at high redshifts, significantly enhancing the constrains on cosmological parameters. The main specific parameter constraints are as follows: for the $\Lambda$CDM model, FAST achieves $\sigma(\Omega_{\mathrm{m}})=0.0076$ and $\sigma(H_{0})=0.51\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$. For the $w$CDM model, the constraint is $\sigma(w)=0.036$. For the $w_{0}w_{a}$ model, the constraints are $\sigma(w_{0})=0.14$, and $\sigma(w_{a})=0.57$. Compared to SKA1-Mid, FAST also has higher sensitivity and lower receiver noise, which enables it to achieve better parameter constraints. For the dark-energy EoS parameters, there is an improvement of approximately 20% for $w$, 12.5% for $w_{0}$, and 21.9% for $w_{a}$.
In the $\Lambda$CDM model, FASTA provides best constraints, $\sigma(\Omega_{m})=3.6\times 10^{-3}$ and $\sigma(H_{0})=0.26\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, which surpasses other instruments and CBS. In the $w$CDM model, FASTA also provides best constraints, $\sigma(\Omega_{\mathrm{m}})=0.0037$, $\sigma(H_{0})=0.39\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, and $\sigma(w)=0.02$. For the parameter $w$, FASTA represents an improvement of approximately 55.56% compared to SKA1-Mid, 44.44% compared to FAST, and 16.67% compared to CBS. In the $w_{0}w_{a}$ model, FASTA achieves highly precise parameter constraints, with $\sigma(\Omega_{\mathrm{m}})=0.012$, $\sigma(H_{0})=0.89\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, $\sigma(w_{0})=0.09$, and $\sigma(w_{a})=0.33$. For $w_{0}$ and $w_{a}$, FASTA represents an improvement of approximately 43.75% and 54.79% compared to SKA1-Mid, and 35.71% and 42.11% compared to FAST, while achieving constraints comparable to CBS using 21 cm IM alone.
III.4 Phased array feed
In this subsection, we discuss the enhancements and challenges of upgrading the feed of FAST to PAF for 21 cm IM. PAF can achieve various beamforming modalities by adjusting the phase and amplitude of each element in the array, thereby generating multiple beams simultaneously. This capability covers a larger field of view ($\mathrm{FOV}$) and enables rapidly beam steering and scanning Landon et al. (2010); Hotan et al. (2021). At the same time, different beamforming strategies significantly impact noise characteristics, and data calibration and processing. We considered two beamforming strategies in terms of survey efficiency.
@@ III.4 Phased array feed
$N_{\mathrm{b}}=\mathrm{FOV}\cdot\left(\frac{\nu D}{c}\right)^{2}.$ (13)
In this case, $C^{\mathrm{N}}$ follows the same pattern as described in Ref. Bull et al. (2015)
$C^{\mathrm{N}}(\bm{q},y)\rightarrow C^{\mathrm{N}}(\bm{q},y)\times\begin{cases}1,&\nu>\nu_{\text{crit }}\\ \left(\nu_{\text{crit }}/\nu\right)^{2},&\nu\leqslant\nu_{\text{crit }}\end{cases}.$ (14)
We assume $\nu_{\mathrm{crit}}=1420\ \mathrm{MHz}$. As shown in Figure 5, the errors increase significantly at low frequencies. This is because at low frequencies, the beam size increases, resulting in a decrease in $N_{\mathrm{b}}$.
Another scenario is a fixed $N_{\mathrm{b}}$. We assume the case of 100 beams. As shown in Figure 5, the increase in the $N_{\mathrm{b}}$ significantly enhances survey efficiency, achieving best constraint result of 2% around $z=1$. However, in practice, the $\mathrm{FOV}$ does not increase indefinitely as the beam size increases at lower frequencies. As the $\mathrm{FOV}$ certainly deviates from the boresight, the survey efficiency decreases. This requires actual measurements to PAF, but here we make an ideal estimate. The real situation is expected to lie between the fixed $\mathrm{FOV}$ and the fixed $N_{\mathrm{b}}$ scenarios.
IV Conclusion
In this paper, we explore several potential upgrades to FAST using 21 cm IM to enhance cosmological research. Those upgrades include FAST with a wide-band receiver, the combination of 40 m dishes and FAST, FASTA, and FAST with a PAF. We simulated the full 21 cm IM power spectrum to constrain the BAO and RSD parameters $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$, and used these results to evaluate the capabilities of these different configurations. Finally, we used the simulated BAO and RSD data to constrain the cosmological parameters of the $\Lambda$CDM, $w$CDM, and $w_{0}w_{a}$ models.

### v2

@@ 2.4 Instrument parameters
2.5 The Fisher matrix
@@
$C^{\mathrm{tot}}=C^{\mathrm{S}}+C^{\mathrm{N}}+C^{\mathrm{F}}.$ (2.9)
For the combination of different telescopes, we consider two scenarios. The first scenario involves two telescopes in close geographic proximity observing the same area of the sky, such as combining 40-m antennas and FAST. This approach maximizes the reduction of thermal noise impact on observations. Since $C^{\mathrm{S}}$ and $C^{\mathrm{F}}$ are only related to the observed sky area, they remain fixed. Different telescopes have varying $C^{\mathrm{N}}$ due to their differing thermal noise integration effects. Therefore, the total covariance matrix can be expressed as
$C^{\mathrm{tot}}=C^{\mathrm{S}}+\left((C_{1}^{\mathrm{N}})^{-1}+(C_{2}^{\mathrm{N}})^{-1}\right)^{-1}+C^{\mathrm{F}},$ (2.10)
@@
$C^{\mathrm{tot}}=\left((C_{1}^{\mathrm{tot}})^{-1}+(C_{2}^{\mathrm{tot}})^{-1}\right)^{-1},$ (2.11)
where $C_{1}^{\mathrm{tot}}$ and $C_{2}^{\mathrm{tot}}$ are the total covariance matrices of the different telescopes.
After obtaining the total covariance matrix $C^{\mathrm{tot}}$, we can transform it into the Fisher matrix of observables $\{p_{i}\}$ [26]
$F_{ij}=\frac{1}{2}U_{\text{bin}}\int\frac{\mathrm{d}^{2}q\,\mathrm{d}y}{(2\pi)^{3}}\left[\frac{\partial\ln C^{\mathrm{tot}}(\bm{q},y)}{\partial p_{i}}\frac{\partial\ln C^{\mathrm{tot}}(\bm{q},y)}{\partial p_{j}}\right].$ (2.12)
In this paper, we mainly calculate the Fisher matrix for several key observables to constain the evolution of LSS, as well as the growth and geometry of the universe. These include the transverse distance indicator $D_{\mathrm{A}}(z)$, radial distance indicator $H(z)$ and the parameter combination $f\sigma_{8}(z)$ from redshift space distortion (RSD), where $f(z)$ is growth rate and $\sigma_{8}$ is the amplitude of matter density fluctuations on a scale of $8h^{-1}$ Mpc (with $h$ the dimensionless Hubble constant).
2.6 Cosmological parameter constraints
We use the Fisher matrix to propagate errors from observational errors to the BAO parameters $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$. Then, we use the errors of BAO parameters to form the likelihood function for cosmological models. Finally, we employ Markov Chain Monte Carlo (MCMC) analysis to constrain the cosmological parameters [96]. Our considered parameter set is $\{H_{0},\Omega_{\mathrm{m}},\Omega_{\mathrm{b}},\sigma_{8},n_{\mathrm{s}},w_{0},w_{a}\}$. We assume a flat universe with $\Omega_{K}=0$ and fix the total neutrino mass $\sum m_{\nu}=\mathrm{0.06\ eV}$. For the free parameters, we assume the priors within the following ranges: $H_{0}\in[50,80]$ km s-1 Mpc-1, $\Omega_{\mathrm{m}}\in[0,0.7]$, $\Omega_{\mathrm{b}}\in[0.03,0.07]$, $\sigma_{8}\in[0,2]$, $n_{\mathrm{s}}\in[0.8,1.2]$, $w_{0}\in[-3,5]$, and $w_{a}\in[-5,5]$. Although BAO alone cannot constrain $n_{\mathrm{s}}$ and $\Omega_{\mathrm{b}}$, we consider them as free parameters to maintain consistency of the parameter space with the cases when considering CMB and supernova (SN) data.
To compare with the mainstream observations, we calculated the constraints from the latest observations. The data used include: CMB – Planck 2018 $-\ell$ TT, TE, EE, and $-\ell$ TT, EE, and PR4 lensing data [97, 98, 99, 100, 101]; BAO – DESI data [2]; and SN – Pantheon+ data [102]. The data combination of CMB, BAO, and SN in this paper is also abbreviated as CBS.
The equation of state (EoS) of dark energy is given by $w=p/\rho$, where $p$ is the pressure and $\rho$ is the energy density. We employ different models to describe its dynamics. (i) $\Lambda$CDM: $w=-1$. (ii) $w$CDM: $w(z)=w$, where $w$ is a constant EoS for dark energy. (iii) $w_{0}w_{a}$: $w$ varies with redshift according to the parameterization $w(z)=w_{0}+w_{a}z/(1+z)$, where $w_{0}$ is the present-day value of the EoS parameter and $w_{a}$ describes its evolutionary behavior.
@@ 2.6 Cosmological parameter constraints
3 Results and discussion
@@
3.1 Combination of FAST and SKA-Mid
It should be noted that for SKA1-Mid, we consider the single dish mode. This is because, in interferometer mode, the maximum observable scale is determined by the shortest baseline. Although SKA1-Mid has an advantage in angular resolution in interferometer mode, it lacks very short baselines, which limits its ability to observe large angular scales. Its sparse array design results in low baseline density and does not achieve good uv coverage, making it unsuitable for 21 cm IM. Improving its performance in interferometer mode would require a denser antenna array, like the CHIME design. In single dish mode, the maximum observable scale is determined by the aperture of the telescope, allowing SKA1-Mid to observe larger scales. Single dish mode also offers other advantages: a wider field of view, more beams, enabling broader sky coverage within limited observation time. Therefore, single dish mode is an optimal choice for 21 cm IM on cosmological scales [26, 37].
From Figure 1, it is evident that FAST and SKA1-Mid have distinct advantages across different redshift ranges. SKA1-Mid achieves tight constraints within the range of $0.80.35<z<$ 0.8 due to its high survey efficiency. In this range, the relative errors in BAO parameters are typically below 3% (see also Ref. [33]). On the other hand, FAST attains robust constraints within the range of $2.50.8<z<$ 2.5 because of its high transverse resolution and sensitivity. In this range, the relative errors in BAO parameters are typically around or below 4%.
Their combined observations enable precise measurement of BAO parameters across the redshift range of $2.50<z<$ 2.5, achieving constraints on $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ to below 2% for most of this range, and to as low as 1% at $z\approx 1$.
3.2 Combination of FAST and 40-m antennas
In this subsection, we discuss the extent to which incorporating 40-m antennas with FAST enhances the constraints on BAO parameters in 21 cm IM. An array of 40-m antennas, located in close geographical proximity to FAST, enhances the survey efficiency by enabling observations of the same sky regions.
Figure 2 illustrates the fractional constraints achievable on $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ using 21 cm IM with an array of 40-m antennas, FAST, and their combined configuration. The 40-m antennas perform better at low redshifts because they have higher survey efficiency. FAST is more effective at higher redshifts due to its greater sensitivity and resolution. The combination of FAST and 40-m antennas significantly reduces observational errors, particularly at high redshifts. The relative errors in BAO parameters are typically below 4%, with errors reaching as low as 2% at $z\approx 1$.
For the configurations of $6\times 40$-m antennas and $24\times 40$-m antennas combined with FAST, Figure 2 shows that the larger array provides only marginal improvements. This result suggests that the combination of FAST with $6\times 40$-m antennas offers a more optimal configuration for the 21 cm IM survey.
Figure 2: Relative errors of $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$ for 21 cm IM with FAST, 40-m antennas, FAST+$6\times 40$-m antennas, and FAST+$24\times 40$-m antennas.
@@ 3.2 Combination of FAST and 40-m antennas
Figure 3: Constraints (68.3% and 95.4% confidence level) on the dark energy models using simulated data of 21 cm IM with SKA1-Mid, FAST, FASTA, and observational data of CBS. Left, middle, and right panels show the constraints on the $\Lambda$CDM, $w$CDM, and $w_{0}w_{a}$ models, respectively.
3.3 FASTA
In this subsection, we present the cosmological parameters constraints by using simulations of 21 cm IM data with different instruments and current observational data from CBS. Figure 3 shows the parameter constraints within the 68.3% and 95.4% confidence intervals for the $\Lambda$CDM, $w$CDM, and $w_{0}w_{a}$ models. The specific values are listed in Table 3.
Compared to the results in Ref. [33] for FAST ($0.350<z<$ 0.35), a wide-band receiver proposed in this paper utilizes the resolution advantage of the FAST at high redshifts, significantly enhancing the constrains on cosmological parameters. The main specific parameter constraints are as follows: for the $\Lambda$CDM model, FAST achieves $\sigma(\Omega_{\mathrm{m}})=0.0076$ and $\sigma(H_{0})=0.51\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$. For the $w$CDM model, the constraint is $\sigma(w)=0.036$. For the $w_{0}w_{a}$ model, the constraints are $\sigma(w_{0})=0.14$, and $\sigma(w_{a})=0.57$. Compared to SKA1-Mid, FAST also has higher sensitivity and lower receiver noise, which enables it to achieve better parameter constraints. For the dark-energy EoS parameters, there is an improvement of approximately 20% for $w$, 12.5% for $w_{0}$, and 21.9% for $w_{a}$.
In the $\Lambda$CDM model, FASTA provides best constraints, $\sigma(\Omega_{m})=3.6\times 10^{-3}$ and $\sigma(H_{0})=0.26\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, which surpasses other instruments and CBS. In the $w$CDM model, FASTA also provides best constraints, $\sigma(\Omega_{\mathrm{m}})=0.0037$, $\sigma(H_{0})=0.39\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, and $\sigma(w)=0.02$. For the parameter $w$, FASTA offers an improvement of approximately 55.56% compared to SKA1-Mid, 44.44% compared to FAST, and 16.67% compared to CBS. In the $w_{0}w_{a}$ model, FASTA achieves highly precise parameter constraints, with $\sigma(\Omega_{\mathrm{m}})=0.012$, $\sigma(H_{0})=0.89\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{~{}Mpc}^{-1}$, $\sigma(w_{0})=0.09$, and $\sigma(w_{a})=0.33$. For $w_{0}$ and $w_{a}$, FASTA provides an improvement of approximately 43.75% and 54.79% compared to SKA1-Mid, and 35.71% and 42.11% compared to FAST, while achieving constraints comparable to CBS using 21 cm IM alone.
@@ 3.3 FASTA
3.4 Phased array feed
In this subsection, we discuss the enhancements and challenges of upgrading the feed of FAST to PAF for 21 cm IM. PAF can achieve various beamforming modalities by adjusting the phase and amplitude of each element in the array, thereby generating multiple beams simultaneously. This capability covers a larger field of view ($\mathrm{FOV}$) and enables rapidly beam steering and scanning [75, 76]. At the same time, different beamforming strategies significantly impact noise characteristics, and data calibration and processing. We considered two beamforming strategies in terms of survey efficiency.
@@ 3.4 Phased array feed
$N_{\mathrm{b}}=\mathrm{FOV}\cdot\left(\frac{\nu D}{c}\right)^{2}.$ (3.1)
In this case, $C^{\mathrm{N}}$ follows the same pattern as described in Ref. [26]
$C^{\mathrm{N}}(\bm{q},y)\rightarrow C^{\mathrm{N}}(\bm{q},y)\times\begin{cases}1,&\nu>\nu_{\text{crit }}\\ \left(\nu_{\text{crit }}/\nu\right)^{2},&\nu\leqslant\nu_{\text{crit }}\end{cases}.$ (3.2)
Here we have $\nu_{\mathrm{crit}}=1420\ \mathrm{MHz}$. As shown in Figure 5, the errors increase significantly at low frequencies. This is because at low frequencies, the beam size increases, resulting in a decrease in $N_{\mathrm{b}}$.
Another scenario is a fixed $N_{\mathrm{b}}$. We assume the case of 100 beams. As shown in Figure 5, the increase in $N_{\mathrm{b}}$ significantly enhances survey efficiency, achieving best constraint result of 2% around $z=1$. However, in practice, the $\mathrm{FOV}$ does not increase endlessly as the beam size increases at lower frequencies. As the $\mathrm{FOV}$ certainly deviates from the boresight, the survey efficiency decreases. This requires actual measurements to PAF, but here we make an ideal estimate. The real situation is expected to lie between the fixed $\mathrm{FOV}$ and the fixed $N_{\mathrm{b}}$ scenarios.
4 Conclusion
In this paper, we explore several potential upgrades to FAST using 21 cm IM to enhance cosmological research. Those upgrades include FAST with a wide-band receiver, the combination of 40-m antennas and FAST, FASTA, and FAST with a PAF. We simulated the full 21 cm IM power spectrum to constrain the BAO and RSD parameters $D_{\mathrm{A}}(z)$, $H(z)$, and $f\sigma_{8}(z)$, and used these results to evaluate the capabilities of these different configurations. Finally, we used the simulated BAO and RSD data to constrain the cosmological parameters of the $\Lambda$CDM, $w$CDM, and $w_{0}w_{a}$ models.